### v1

%-------------------------------------------% ̵Paper ̵Head%-------------------------------------------“title–Recursive ̵Interferometric ̵Surface-wave ̵Suppression ̵For ̵Improved ̵Reflection ̵Imaging˝“author[1,*]–Faezeh ̵Shirmohammadi˝“author[1]–Deyan ̵Draganov˝“author[1]–Ranajit ̵Ghose˝“author[1]–Eric ̵Verschuur˝“author[1]–Kees ̵Wapenaar˝“affil[1]–“small ̵Department ̵of ̵Geoscience ̵and ̵Engineering, ̵Delft ̵University ̵of ̵Technology, ̵2628 ̵CN ̵Delft, ̵The ̵Netherlands˝“affil[*]–Corresponding ̵author: ̵“texttt–f.shirmohammadi@tudelft.nl˝˝“renewcommand“Authands– ̵and ̵˝“date–˝ ̵ ̵% ̵Remove ̵date“begin–document˝“maketitle“begin–abstract˝“onehalfspacingHigh-resolution ̵seismic ̵reflections ̵are ̵essential ̵for ̵imaging ̵and ̵monitoring ̵applications. ̵In ̵seismic ̵land ̵surveys ̵using ̵sources ̵and ̵receivers ̵at ̵the ̵surface, ̵surface ̵waves ̵often ̵dominate, ̵masking ̵the ̵reflections.In ̵this ̵study, ̵we ̵demonstrate ̵the ̵efficacy ̵of ̵a ̵two-step ̵procedure ̵to ̵suppress ̵surface ̵waves ̵in ̵an ̵active-source ̵reflection ̵seismic ̵dataset. ̵First, ̵we ̵apply ̵seismic ̵interferometry ̵(SI) ̵by ̵cross-correlation, ̵turning ̵receivers ̵into ̵virtual ̵sources ̵to ̵estimate ̵the ̵dominant ̵surface ̵waves. ̵Then, ̵we ̵perform ̵adaptive ̵subtraction ̵to ̵minimise ̵the ̵difference ̵between ̵the ̵surface ̵waves ̵in ̵the ̵original ̵data ̵and ̵the ̵result ̵of ̵SI. ̵We ̵propose ̵a ̵new ̵approach ̵where ̵the ̵initial ̵suppression ̵results ̵are ̵used ̵for ̵further ̵iterations, ̵followed ̵by ̵adaptive ̵subtraction. ̵This ̵technique ̵aims ̵to ̵enhance ̵the ̵efficacy ̵of ̵data-driven ̵surface-wave ̵suppression ̵through ̵an ̵iterative ̵process.We ̵use ̵a ̵2D ̵seismic ̵reflection ̵dataset ̵from ̵Scheemda, ̵situated ̵in ̵the ̵Groningen ̵province ̵of ̵the ̵Netherlands, ̵to ̵illustrate ̵the ̵technique's ̵efficiency. ̵A ̵comparison ̵between ̵the ̵data ̵after ̵recursive ̵interferometric ̵surface-wave ̵suppression ̵and ̵the ̵original ̵data ̵across ̵time ̵and ̵frequency-wavenumber ̵domains ̵shows ̵significant ̵suppression ̵of ̵the ̵surface ̵waves, ̵enhancing ̵visualization ̵of ̵the ̵reflections ̵for ̵following ̵subsurface ̵imaging ̵and ̵monitoring ̵studies. ̵“par“vspace–“baselineskip˝“noindent ̵“textbf–Keywords˝: ̵Data ̵processing, ̵Seismic ̵Interferometry, ̵Seismics, ̵Surface ̵waves, ̵Groningen, ̵Reflections.“end–abstract˝“newpage%-------------------------------------------% ̵Paper ̵Body%-------------------------------------------%--- ̵Section ̵---%“section–Introduction˝Seismic ̵data ̵acquired ̵on ̵land ̵is ̵often ̵contaminated ̵by ̵surface ̵waves, ̵which ̵are ̵a ̵significant ̵noise ̵source ̵for ̵reflection ̵seismic ̵studies. ̵The ̵surface ̵waves ̵often ̵have ̵a ̵velocity ̵and ̵frequency ̵content ̵similar ̵to ̵those ̵of ̵the ̵investigated ̵reflections, ̵which ̵makes ̵it ̵challenging ̵to ̵suppress ̵them ̵from ̵the ̵dataset.Conventionally, ̵surface ̵waves ̵are ̵suppressed ̵during ̵data ̵processing ̵using ̵methods ̵such ̵as ̵frequency-offset ̵(f-x) ̵(“cite–Yilmaz2001˝), ̵frequency-wavenumber ̵(f-k), ̵or ̵bandpass ̵filtering. ̵However, ̵these ̵approaches ̵can ̵prove ̵ineffective ̵when ̵surface ̵waves ̵are ̵scattered ̵and ̵overlap ̵with ̵the ̵frequency ̵and ̵moveout ̵of ̵the ̵reflected ̵body ̵waves ̵that ̵we ̵intend ̵to ̵preserve. ̵The ̵f-k ̵filter ̵may ̵result ̵in ̵artefacts ̵due ̵to ̵signal ̵distortion ̵and ̵spatial ̵correlation ̵of ̵the ̵background ̵noise ̵because ̵the ̵surface-wave ̵energy ̵may ̵be ̵distributed ̵over ̵a ̵significant ̵range ̵of ̵the ̵f-k ̵spectrum ̵(“cite–KONSTANTAKI201528˝), ̵thus ̵further ̵lowering ̵the ̵quality ̵of ̵the ̵reflections. ̵“parRecently, ̵the ̵prediction ̵of ̵surface ̵waves ̵with ̵seismic ̵interferometry ̵(SI) ̵and ̵their ̵subsequent ̵adaptive ̵subtraction ̵from ̵seismic ̵reflection ̵data ̵has ̵emerged ̵as ̵a ̵technique ̵for ̵suppressing ̵surface ̵waves ̵(“cite–Dong2006˝; ̵“cite–Halliday2010˝). ̵In ̵SI, ̵seismic ̵observations ̵from ̵various ̵receiver ̵locations ̵are, ̵for ̵example, ̵cross-correlated ̵to ̵retrieve ̵new ̵seismic ̵responses ̵from ̵virtual ̵sources ̵positioned ̵at ̵the ̵receivers' ̵locations. ̵This ̵process ̵enables ̵the ̵retrieval ̵of ̵seismic ̵responses ̵between ̵pairs ̵of ̵receivers. ̵For ̵suppression, ̵the ̵retrieved ̵responses ̵are ̵then ̵subtracted ̵from ̵the ̵original ̵field ̵recordings ̵using ̵least-squares ̵matching, ̵resulting ̵in ̵data ̵with ̵suppressed ̵surface ̵waves. ̵This ̵suppression ̵technique ̵is ̵usually ̵called ̵interferometric ̵surface-wave ̵suppression. ̵“parIn ̵previous ̵studies, ̵the ̵interferometric ̵surface-wave ̵suppression ̵was ̵applied ̵in ̵a ̵non-recursive ̵way ̵on ̵the ̵data. ̵“textcite–Halliday2010˝ ̵demonstrated ̵its ̵effectiveness ̵in ̵the ̵context ̵of ̵hydrocarbon ̵exploration, ̵while ̵“textcite–KONSTANTAKI201528˝ ̵and ̵“textcite–Liu2018SeismicWaves˝ ̵showcased ̵its ̵utility ̵for ̵near-surface ̵applications. ̵Moreover, ̵“textcite–Balestrini2020ImprovedSuppression˝ ̵demonstrated ̵its ̵application ̵for ̵deep ̵mineral ̵explorations. ̵Here, ̵we ̵suggest ̵a ̵new ̵approach ̵using ̵the ̵first ̵output ̵of ̵the ̵interferometric ̵surface-wave ̵suppression ̵for ̵more ̵iterations. ̵We ̵term ̵this ̵technique ̵"Recursive ̵Interferometric ̵Surface-wave ̵Suppression" ̵(RISS). ̵This ̵technique ̵aims ̵to ̵enhance ̵the ̵efficacy ̵of ̵the ̵data-driven ̵surface-wave ̵suppression ̵through ̵an ̵iterative ̵procedures. ̵“parIn ̵this ̵study, ̵we ̵demonstrate ̵RISS ̵on ̵a ̵2D ̵reflection ̵dataset ̵acquired ̵in ̵Scheemda, ̵Groningen ̵province, ̵the ̵Netherlands. ̵By ̵using ̵RISS, ̵we ̵aim ̵to ̵enhance ̵the ̵visualisation ̵of ̵reflections, ̵which ̵can ̵provide ̵clearer ̵images ̵of ̵the ̵subsurface ̵structures ̵and ̵enhance ̵the ̵overall ̵interpretation ̵of ̵the ̵seismic ̵data. ̵Such ̵advancements ̵are ̵particularly ̵critical ̵for ̵Groningen, ̵where ̵gas ̵production ̵has ̵resulted ̵in ̵induced ̵seismicity ̵since ̵1963 ̵(“cite–Muntendam-Bos2022AnNetherlands˝). ̵We ̵evaluate ̵the ̵RISS ̵results ̵in ̵comparison ̵with ̵those ̵from ̵other ̵techniques ̵such ̵as ̵time ̵muting ̵and ̵f-k ̵filtering.“parBelow, ̵we ̵first ̵present ̵in ̵Section ̵2 ̵the ̵methodology ̵of ̵RISS. ̵This ̵will ̵be ̵followed ̵by ̵a ̵description ̵of ̵the ̵seismic ̵data ̵acquisition ̵in ̵Section ̵3, ̵the ̵results ̵in ̵Section ̵4, ̵and ̵then ̵a ̵discussion ̵and ̵the ̵conclusions. ̵“par%--- ̵Section ̵---%“section–Methodology˝“label–Method˝In ̵our ̵proposed ̵approach, ̵SI ̵is ̵employed ̵first ̵to ̵retrieve ̵the ̵dominant ̵surface ̵waves. ̵The ̵retrieved ̵surface-wave ̵energy ̵is ̵subsequently ̵adaptively ̵subtracted ̵from ̵the ̵dataset. ̵Following ̵this, ̵the ̵obtained ̵data ̵is ̵utilised ̵to ̵iterate ̵through ̵these ̵two ̵steps, ̵contributing ̵to ̵the ̵improvement ̵of ̵the ̵reflection ̵resolution. ̵This ̵section ̵outlines ̵the ̵implementation ̵of ̵RISS.%-------------------------------------------“subsection–Surface-wave ̵retrieval ̵by ̵seismic ̵interferometry˝SI ̵refers ̵to ̵the ̵method ̵of ̵retrieving ̵new ̵seismic ̵responses, ̵for ̵example ̵between ̵two ̵receivers, ̵using ̵most ̵commonly ̵cross-correlation, ̵as ̵if ̵there ̵were ̵a ̵source ̵at ̵one ̵of ̵the ̵receiver ̵locations ̵(“cite–Wapenaar2006GreensInterferometry˝; ̵“cite–Wapenaar2010˝). ̵In ̵an ̵active-source ̵survey, ̵this ̵process ̵is ̵usually ̵achieved ̵by ̵cross-correlating ̵the ̵recordings ̵at ̵the ̵two ̵receivers ̵and ̵then ̵stacking ̵the ̵virtual ̵traces ̵over ̵all ̵available ̵active ̵sources ̵(“cite–Halliday2008˝). ̵So, ̵the ̵retrieved ̵response ̵between ̵two ̵receivers ̵at ̵positions ̵\$“mathbf–x˝˙–A˝\$ ̵and ̵\$“mathbf–x˝˙–B˝\$ ̵can ̵be ̵expressed ̵in ̵the ̵frequency ̵domain, ̵as ̵described ̵by ̵“textcite–Halliday2007InterferometricRemoval˝:%%%%%%%%%%%%%% ̵eq ̵5˙1 ̵%%%%%%%%%%%%%%“begin–equation˝ ̵ ̵ ̵%“intertext–where˝“, ̵ ̵ ̵“hat–G˝^*(“mathbf–x˝˙–B˝,“mathbf–x˝˙–A˝,“omega) ̵+ ̵“hat–G˝(“mathbf–x˝˙–B˝,“mathbf–x˝˙–A˝,“omega) ̵“propto ̵“sum˙–n=1˝^–N˝ ̵ ̵ ̵“hat–G˝^*(“mathbf–x˝˙–A˝,x˙–n˝,“omega) ̵ ̵“hat–G˝(“mathbf–x˝˙–B˝,x˙–n˝,“omega) ̵, ̵ ̵ ̵“label–eq:5˙1˝“end–equation˝“noindent ̵where ̵\$“hat–G˝(“mathbf–x˝˙–B˝,x˙–n˝,“omega)\$ ̵is ̵the ̵frequency-domain ̵response ̵of ̵a ̵recording ̵at ̵receiver ̵\$“mathbf–x˝˙–B˝\$ ̵and ̵\$“hat–G˝(“mathbf–x˝˙–A˝,x˙–n˝,“omega)\$ ̵is ̵the ̵frequency-domain ̵response ̵of ̵a ̵recording ̵at ̵receiver ̵\$“mathbf–x˝˙–A˝\$, ̵both ̵from ̵a ̵source ̵at ̵\$x˙–n˝\$ ̵at ̵the ̵Earth's ̵surface. ̵The ̵left-hand ̵side ̵of ̵the ̵equation ̵represents ̵the ̵response ̵between ̵the ̵two ̵receivers ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵and ̵\$“mathbf–x˝˙–B˝\$ ̵at ̵the ̵surface, ̵implying ̵that ̵we ̵turn ̵the ̵receiver ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵into ̵a ̵virtual ̵source. ̵The ̵symbol ̵(\$*\$) ̵denotes ̵the ̵complex ̵conjugate ̵in ̵the ̵frequency ̵domain, ̵while ̵\$N\$ ̵represents ̵the ̵total ̵number ̵of ̵active ̵sources ̵at ̵the ̵surface.“parIn ̵a ̵laterally ̵homogeneous ̵2D ̵medium, ̵sources ̵at ̵points ̵in ̵line ̵with ̵the ̵receivers ̵contribute ̵to ̵the ̵retrieval ̵of ̵direct ̵surface-wave ̵arrivals ̵since ̵they ̵are ̵all ̵in ̵the ̵stationary-phase ̵region, ̵but ̵only ̵a ̵few ̵points ̵are ̵stationary ̵to ̵retrieve ̵pseudo-physical ̵(“cite–Loer2013DiffractedSRI˝) ̵or ̵ghost ̵reflections ̵(–“cite–Draganov2012MonitoringInterferometry˝; ̵“cite–shirmohamamdigroningen˝) ̵and ̵scattered ̵surface ̵waves. ̵This ̵way, ̵the ̵results ̵retrieved ̵by ̵SI ̵will ̵be ̵dominated ̵by ̵surface ̵waves ̵(“cite–Balestrini2020ImprovedSuppression˝).“parFigure~“ref–fig:5˙1˝ ̵shows ̵a ̵schematic ̵representation ̵of ̵SI ̵for ̵retrieving ̵direct ̵arrivals, ̵including ̵surface ̵waves. ̵By ̵correlating ̵the ̵recording ̵at ̵\$“mathbf–x˝˙–B˝\$ ̵from ̵the ̵active ̵source ̵at ̵\$“mathbf–x˝\$ ̵in ̵Figure~“ref–fig:5˙1˝a ̵with ̵a ̵recording ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵in ̵Figure~“ref–fig:5˙1˝c, ̵the ̵virtual ̵response ̵between ̵\$“mathbf–x˝˙–B˝\$ ̵and ̵\$“mathbf–x˝˙–A˝\$ ̵is ̵retrieved, ̵as ̵illustrated ̵by ̵the ̵purple ̵arrow ̵in ̵Figure~“ref–fig:5˙1˝d ̵at ̵causal ̵times ̵(the ̵causal ̵part ̵refers ̵to ̵times ̵later ̵than ̵the ̵zero ̵time). ̵Similarly, ̵the ̵virtual ̵response ̵between ̵another ̵receiver ̵at ̵\$“mathbf–x˝˙–B'˝\$ ̵and ̵a ̵receiver ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵is ̵retrieved ̵by ̵correlating ̵the ̵response ̵at ̵\$“mathbf–x˝˙–B'˝\$ ̵in ̵Figure~“ref–fig:5˙1˝b ̵with ̵\$“mathbf–x˝˙–A˝\$ ̵in ̵Figure~“ref–fig:5˙1˝c, ̵as ̵depicted ̵by ̵the ̵orange ̵arrow ̵in ̵Figure~“ref–fig:5˙1˝d ̵at ̵acausal ̵times ̵(the ̵acausal ̵part ̵refers ̵to ̵times ̵earlier ̵than ̵the ̵zero ̵time). ̵In ̵both ̵cases, ̵the ̵receiver ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵acts ̵as ̵a ̵virtual ̵source ̵as ̵shown ̵by ̵the ̵blue ̵explosion ̵in ̵Figure~“ref–fig:5˙1˝d. ̵We ̵repeat ̵this ̵procedure ̵for ̵all ̵active ̵sources, ̵e.g., ̵as ̵shown ̵in ̵Figure~“ref–fig:5˙1˝e ̵for ̵another ̵active ̵source ̵at ̵\$“mathbf–x˝^–'˝\$. ̵Finally, ̵the ̵Green's ̵function ̵and ̵its ̵time-reversal ̵between ̵the ̵virtual ̵source ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵and ̵other ̵receivers ̵at ̵\$“mathbf–x˝˙–B˝\$, ̵\$“mathbf–x˝˙–B'˝\$, ̵and ̵ ̵\$“mathbf–x˝˙–B''˝\$ ̵are ̵retrieved ̵by ̵stacking ̵all ̵virtual ̵responses ̵such ̵as ̵those ̵shown ̵in ̵Figures~“ref–fig:5˙1˝d ̵and~“ref–fig:5˙1˝e. ̵“parWhen ̵we ̵want ̵to ̵apply ̵this ̵technique ̵to ̵a ̵field ̵dataset, ̵there ̵are ̵certain ̵issues ̵that ̵need ̵to ̵be ̵addressed ̵in ̵order ̵to ̵improve ̵the ̵resolution ̵of ̵the ̵retrieved ̵responses. ̵“parFirst, ̵we ̵aim ̵to ̵retrieve ̵the ̵direct ̵surface ̵waves ̵with ̵SI. ̵So, ̵it ̵is ̵required ̵that ̵all ̵receivers ̵be ̵considered ̵on ̵the ̵same ̵side ̵of ̵the ̵active ̵source, ̵e.g, ̵for ̵an ̵active ̵source ̵at ̵\$“mathbf–x˝\$, ̵we ̵correlate ̵the ̵response ̵for ̵receivers ̵\$“mathbf–x˝˙–B˝\$ ̵and ̵\$“mathbf–x˝˙–B'˝\$ ̵located ̵on ̵the ̵same ̵side ̵as ̵the ̵virtual ̵source ̵at ̵\$“mathbf–x˝˙–A˝\$, ̵as ̵shown ̵in ̵Figures~“ref–fig:5˙1˝a, ̵b, ̵and ̵c. ̵In ̵the ̵same ̵way, ̵for ̵an ̵active ̵source ̵at ̵\$“mathbf–x˝^–'˝\$, ̵we ̵consider ̵all ̵receivers ̵because ̵they ̵are ̵on ̵the ̵same ̵side ̵as ̵the ̵virtual ̵source ̵at ̵\$“mathbf–x˝˙–A˝\$, ̵as ̵shown ̵in ̵Figure~“ref–fig:5˙1˝e. ̵This ̵condition ̵is ̵dictated ̵by ̵the ̵theory ̵of ̵SI ̵by ̵cross-correlation, ̵which ̵states ̵that ̵the ̵sources ̵should ̵surround ̵the ̵receivers, ̵i.e., ̵there ̵must ̵not ̵be ̵sources ̵between ̵the ̵receivers ̵involved ̵in ̵the ̵correlation ̵process. ̵“parSecond, ̵in ̵the ̵case ̵of ̵isotropic ̵illumination ̵of ̵the ̵receivers, ̵a ̵time-symmetric ̵response ̵between ̵the ̵receivers ̵is ̵obtained, ̵as ̵shown ̵in ̵equation~“ref–eq:5˙1˝. ̵Consequently, ̵one ̵could ̵sum ̵the ̵causal ̵and ̵the ̵time-reversed ̵acausal ̵parts ̵of ̵the ̵correlated ̵panels ̵to ̵improve ̵the ̵signal-to-noise ̵ratio. ̵However, ̵in ̵practical ̵situations, ̵when ̵the ̵illumination ̵is ̵not ̵homogeneous ̵from ̵all ̵sides ̵for ̵each ̵pair ̵of ̵receivers, ̵then ̵parts ̵of ̵the ̵response ̵can ̵be ̵retrieved ̵at ̵acausal ̵times ̵and ̵other ̵parts ̵at ̵causal ̵times. ̵Therefore, ̵to ̵enhance ̵the ̵quality ̵of ̵our ̵retrieved ̵responses, ̵we ̵meticulously ̵assess ̵the ̵positions ̵of ̵virtual ̵source-receiver ̵pairs ̵and ̵active ̵sources. ̵Subsequently, ̵we ̵opt ̵to ̵select ̵either ̵the ̵causal ̵or ̵time-reversed ̵acausal ̵part ̵of ̵the ̵correlation ̵panel. ̵“parConsidering ̵the ̵conditions ̵of ̵one-sided ̵distribution ̵of ̵receivers ̵and ̵causality, ̵we ̵limit ̵ourselves ̵to ̵a ̵minimum ̵number ̵of ̵traces ̵for ̵stacking. ̵To ̵maintain ̵a ̵high ̵signal-to-noise ̵ratio, ̵we ̵stack ̵traces ̵only ̵when ̵we ̵have ̵at ̵least ̵half ̵the ̵number ̵of ̵all ̵active ̵sources. ̵We ̵can ̵summarise ̵this ̵as ̵follows:“begin–equation*˝ ̵ ̵ ̵ ̵trace ̵ ̵ ̵ ̵“left“– ̵ ̵ ̵ ̵“begin–array˝–ll˝ ̵ ̵ ̵ ̵ ̵ ̵ ̵“text–causal ̵part, ̵˝ ̵“text–if ̵˝ ̵ ̵m>N/2 ̵ ̵“text– ̵with ̵˝“mathbf–x˝ ̵< ̵“mathbf–x˝˙–A˝ ̵< ̵“mathbf–x˝˙–B˝ ̵“text–, ̵˝“text–or ̵˝ ̵“mathbf–x˝ ̵> ̵“mathbf–x˝˙–A˝ ̵> ̵“mathbf–x˝˙–B˝, ̵“text–or ̵˝ ̵“mathbf–x˝˙–A˝ ̵= ̵“mathbf–x˝˙–B˝ ̵““ ̵ ̵ ̵ ̵ ̵ ̵ ̵ ̵“text–time-reversed ̵acausal, ̵˝ ̵ ̵ ̵“text–if ̵˝ ̵m>N/2 ̵ ̵“text– ̵with ̵˝ ̵“mathbf–x˝ ̵< ̵“mathbf–x˝˙–B˝ ̵< ̵“mathbf–x˝˙–A˝ ̵“text–, ̵˝“text–or ̵˝ ̵“mathbf–x˝ ̵> ̵“mathbf–x˝˙–B˝ ̵> ̵“mathbf–x˝˙–A˝. ̵ ̵ ̵ ̵“end–array˝ ̵ ̵ ̵ ̵“right.“end–equation*˝Here, ̵\$“mathbf–x˝\$ ̵is ̵the ̵position ̵of ̵the ̵active ̵sources, ̵\$“mathbf–x˝˙–A˝\$ ̵is ̵the ̵position ̵of ̵the ̵virtual ̵source, ̵\$“mathbf–x˝˙–B˝\$ ̵is ̵the ̵position ̵of ̵the ̵receiver, ̵and ̵\$m\$ ̵is ̵a ̵number ̵of ̵active ̵sources ̵for ̵stacking ̵which ̵should ̵be ̵higher ̵than ̵half ̵the ̵total ̵number ̵of ̵active ̵sources ̵(\$N\$). ̵For ̵instance, ̵as ̵illustrated ̵in ̵Figure~“ref–fig:5˙1˝d, ̵for ̵the ̵receiver ̵at ̵\$“mathbf–x˝˙–B'˝\$ ̵on ̵the ̵left ̵side ̵of ̵the ̵virtual ̵source ̵at ̵\$“mathbf–x˝˙–A˝\$, ̵we ̵consider ̵the ̵acausal ̵part, ̵as ̵indicated ̵by ̵the ̵orange ̵arrow. ̵Similarly, ̵for ̵a ̵receiver ̵at ̵\$“mathbf–x˝˙–B˝\$, ̵we ̵consider ̵the ̵causal ̵part, ̵indicated ̵by ̵the ̵purple ̵arrow. ̵Furthermore, ̵as ̵depicted ̵in ̵Figure~“ref–fig:5˙1˝e, ̵for ̵the ̵receiver ̵at ̵\$“mathbf–x˝˙–B˝\$ ̵on ̵the ̵right ̵side ̵of ̵the ̵virtual ̵source ̵at ̵\$“mathbf–x˝˙–A˝\$, ̵we ̵consider ̵the ̵acausal ̵part, ̵and ̵for ̵other ̵receivers ̵at ̵\$“mathbf–x˝˙–B'˝\$ ̵and ̵\$“mathbf–x˝˙–B''˝\$, ̵we ̵consider ̵the ̵causal ̵part, ̵as ̵shown ̵by ̵the ̵purple ̵arrows. ̵Considering ̵the ̵above ̵factors, ̵we ̵retrieve ̵the ̵virtual ̵common-source ̵gather ̵for ̵all ̵receivers. ̵“par%%%%%%%%%%%%%%% ̵Figure ̵5-1 ̵%%%%%%%%%%%%%%“begin–figure˝[ht] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.7“textwidth]–Figures/Figure5˙1.png˝ ̵ ̵ ̵ ̵“caption–Schematic ̵representation ̵of ̵seismic ̵interferometry ̵for ̵retrieving ̵surface ̵waves. ̵(a), ̵(b), ̵and ̵(c) ̵The ̵surface ̵wave ̵recorded ̵at ̵\$“mathbf–x˝˙–B˝\$, ̵\$“mathbf–x˝˙–B'˝\$, ̵and ̵\$“mathbf–x˝˙–A˝\$ ̵from ̵the ̵active ̵source ̵at ̵\$“mathbf–x˝\$, ̵respectively. ̵(d) ̵The ̵result ̵of ̵correlating ̵the ̵response ̵at ̵\$“mathbf–x˝˙–A˝\$ ̵with ̵other ̵receivers, ̵and ̵thus ̵turning ̵\$“mathbf–x˝˙–A˝\$ ̵into ̵a ̵virtual ̵source. ̵(e) ̵Same ̵as ̵(d) ̵but ̵for ̵the ̵active ̵source ̵at ̵\$“mathbf–x˝^–'˝\$. ̵The ̵black ̵arrows ̵indicate ̵the ̵surface ̵waves, ̵while ̵the ̵orange ̵and ̵purple ̵arrows ̵represent ̵the ̵outcomes ̵of ̵correlation, ̵considering ̵the ̵causal ̵and ̵acausal ̵parts, ̵respectively.˝ ̵ ̵ ̵ ̵“label–fig:5˙1˝“end–figure˝“par%-------------------------------------------“subsection–Adaptive ̵subtraction˝When ̵each ̵source ̵position ̵in ̵an ̵active-source ̵survey ̵is ̵in ̵close ̵proximity ̵to ̵a ̵receiver ̵position, ̵we ̵are ̵able ̵to ̵identify ̵a ̵corresponding ̵retrieved ̵virtual ̵common-source ̵gather ̵with ̵estimated ̵dominant ̵surface ̵waves ̵for ̵each ̵active ̵source–virtual ̵source ̵pair. ̵These ̵estimates ̵can ̵then ̵be ̵adaptively ̵subtracted ̵from ̵the ̵complete ̵responses ̵of ̵the ̵active ̵source ̵pair ̵(“cite–Halliday2008˝, ̵“cite–Halliday2010˝, ̵“cite–KONSTANTAKI201528˝). ̵“parFigure~“ref–fig:5˙2˝ ̵shows ̵the ̵basic ̵principle ̵of ̵adaptive ̵subtraction. ̵Figure~“ref–fig:5˙2˝a ̵illustrates ̵schematically ̵a ̵simple ̵seismic ̵trace ̵that ̵consists ̵of ̵a ̵primary ̵reflection ̵at ̵100 ̵ms, ̵shown ̵in ̵black, ̵and ̵surface ̵waves ̵at ̵300 ̵ms, ̵350 ̵ms, ̵and ̵400 ̵ms ̵in ̵blue. ̵Figure~“ref–fig:5˙2˝b ̵shows ̵the ̵result ̵of ̵SI, ̵which ̵contains ̵surface ̵waves. ̵By ̵minimising ̵the ̵difference ̵between ̵Figure~“ref–fig:5˙2˝b ̵and ̵Figure~“ref–fig:5˙2˝a, ̵the ̵surface ̵waves ̵in ̵Figure~“ref–fig:5˙2˝a ̵can ̵be ̵suppressed ̵partly, ̵as ̵shown ̵in ̵Figure~“ref–fig:5˙2˝c. ̵The ̵minimisation ̵is ̵done ̵by ̵estimating ̵ ̵a ̵shaping ̵filter ̵“textbf–f˝ ̵that ̵can ̵minimise ̵the ̵following ̵objective ̵function ̵(“cite–Liu2018SeismicWaves˝; ̵“cite–Balestrini2020ImprovedSuppression˝):%%%%%%%%%%%%%% ̵eq ̵5˙2 ̵%%%%%%%%%%%%%%“begin–equation˝ ̵ ̵ ̵“displaystyle ̵“min˙–“boldsymbol–f˝˝ ̵“lVert ̵–D˝ ̵- ̵“boldsymbol–f˝–D˝˙–“text–SW˝˝“rVert^2 ̵, ̵ ̵ ̵“label–eq:5˙2˝“end–equation˝“noindent ̵where ̵\$D\$ ̵is ̵the ̵field ̵dataset ̵with ̵surface ̵waves ̵and ̵\$–D˝˙–SW˝\$ ̵are ̵the ̵surface ̵waves ̵retrieved ̵by ̵SI ̵using ̵the ̵field ̵dataset. ̵The ̵vertical ̵double ̵bars ̵\$——.——^2\$ ̵represent ̵the ̵L2 ̵norm. ̵Equation~“ref–eq:5˙2˝ ̵is ̵solved ̵using ̵an ̵iterative ̵least-squares ̵fit ̵(“cite–Verschuur1992˝). ̵More ̵details ̵can ̵be ̵found ̵in ̵“textcite–Alai˝ ̵and ̵“textcite–seismic-multiple-removal˝. ̵The ̵multiplication ̵between ̵the ̵“textbf–f˝ ̵and ̵\$–D˝˙–SW˝\$ ̵is ̵directly ̵subtracted ̵from ̵\$D\$, ̵giving ̵\$–D˙–ref˝˝\$ ̵which ̵represents ̵the ̵data ̵after ̵surface-wave ̵suppression ̵as ̵“par%%%%%%%%%%%%%% ̵eq ̵5˙3 ̵%%%%%%%%%%%%%%“begin–equation˝ ̵ ̵ ̵–D˙–“text–ref˝˝˝ ̵= ̵–D˝ ̵- ̵“boldsymbol–f˝–D˝˙–“text–SW˝˝. ̵ ̵ ̵“label–eq:5˙3˝“end–equation˝As ̵illustrated ̵in ̵Figure~“ref–fig:5˙2˝c, ̵the ̵data ̵after ̵adaptive ̵subtraction ̵may ̵still ̵contain ̵surface ̵waves ̵due ̵to, ̵for ̵example, ̵errors ̵in ̵estimating ̵higher ̵modes ̵of ̵surface ̵waves. ̵Therefore, ̵we ̵suggest ̵repeating ̵the ̵same ̵step ̵of ̵SI ̵and ̵adaptive ̵subtraction ̵but ̵now ̵using ̵the ̵output ̵of ̵the ̵first ̵adaptive ̵subtraction, ̵represented ̵here ̵as ̵Figure~“ref–fig:5˙2˝c, ̵as ̵input ̵for ̵SI. ̵We ̵estimate ̵surface ̵waves ̵from ̵SI, ̵as ̵shown ̵in ̵Figure~“ref–fig:5˙2˝d, ̵and ̵then ̵adaptively ̵subtract ̵them ̵from ̵Figure~“ref–fig:5˙2˝c. ̵Figure~“ref–fig:5˙2˝e ̵shows ̵the ̵final ̵result, ̵whereby ̵repeating ̵these ̵steps, ̵we ̵increase ̵our ̵chances ̵of ̵suppressing ̵all ̵the ̵surface-wave ̵energy ̵and ̵preserving ̵weak ̵reflections ̵(the ̵black ̵event). ̵We ̵call ̵this ̵technique ̵RISS. ̵Note, ̵that ̵RISS ̵can ̵be ̵applied ̵for ̵one ̵iteration ̵or ̵multiple ̵iterations.%%%%%%%%%%%%%%% ̵Figure ̵5-2 ̵%%%%%%%%%%%%%%“begin–figure˝[htbp] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=1.0“textwidth]–Figures/Figure5˙2.png˝ ̵ ̵ ̵ ̵“caption–Schematic ̵representation ̵of ̵Recursive ̵Interferometric ̵Surface-wave ̵Suppression ̵(RISS). ̵(a) ̵A ̵seismic ̵trace ̵\$D\$ ̵with ̵a ̵primary ̵reflection ̵in ̵black ̵and ̵surface ̵waves ̵in ̵blue, ̵(b) ̵\$–D˝˙–SW˝^1\$ ̵is ̵a ̵result ̵of ̵SI, ̵(c) ̵\$–D˙–ref˝^1˝\$ ̵is ̵the ̵trace ̵after ̵the ̵first ̵iteration ̵of ̵RISS, ̵(d) ̵\$–D˝˙–SW˝^2\$ ̵is ̵a ̵result ̵of ̵SI ̵when ̵using ̵ ̵\$–D˙–ref˝^1˝\$ ̵in ̵(c) ̵as ̵input, ̵(e) ̵\$–D˙–ref˝^2˝\$ ̵after ̵a ̵second ̵iteration ̵of ̵RISS.˝ ̵ ̵ ̵ ̵“label–fig:5˙2˝“end–figure˝“newpage%-------------------------------------------“section–Seismic ̵data ̵acquisition˝We ̵acquired ̵a ̵2D ̵seismic ̵reflection ̵data ̵close ̵to ̵the ̵town ̵of ̵Scheemda ̵in ̵the ̵Groningen ̵province ̵of ̵the ̵Netherlands ̵in ̵2022. ̵Figures~“ref–fig:5˙3˝a ̵and ̵~“ref–fig:5˙3˝b ̵show ̵the ̵location ̵of ̵the ̵site ̵and ̵the ̵geometry ̵of ̵the ̵reflection ̵line, ̵respectively. ̵We ̵employed ̵an ̵electric ̵linear ̵motor ̵system ̵(LMS) ̵seismic ̵vibrator ̵(“cite–Noorlandt2015AMotors˝) ̵as ̵a ̵source, ̵with ̵a ̵spacing ̵of ̵2 ̵m ̵(red ̵stars ̵in ̵Figure~“ref–fig:5˙3˝b), ̵and ̵601 ̵three-component ̵geophone ̵nodes ̵as ̵receivers ̵(the ̵circles ̵in ̵Figure~“ref–fig:5˙3˝b), ̵with ̵a ̵spacing ̵of ̵1 ̵m. ̵The ̵acquisition ̵parameters ̵are ̵summarised ̵in ̵Table~“ref–tab:5˙1˝.“parFor ̵this ̵survey, ̵we ̵used ̵the ̵electrical ̵vibrator ̵in ̵the ̵S-wave ̵mode ̵and ̵oriented ̵in ̵the ̵crossline ̵direction. ̵We ̵then ̵used ̵the ̵data ̵recorded ̵by ̵the ̵crossline ̵horizontal ̵component ̵of ̵the ̵geophones. ̵Because ̵of ̵the ̵orientation ̵of ̵the ̵sources ̵and ̵the ̵receivers, ̵and ̵assuming ̵no ̵scattering ̵from ̵the ̵crossline ̵direction, ̵the ̵horizontally ̵polarised ̵S-waves ̵(SH-waves) ̵we ̵record ̵are ̵generally ̵decoupled ̵from ̵the ̵compressional ̵and ̵vertically ̵polarised ̵S-waves.“par“par“par“par“begin–table˝[htbp] ̵ ̵“caption–Acquisition ̵parameters˝ ̵ ̵“small ̵ ̵“centering ̵ ̵ ̵ ̵“begin–tabular˝–—c—c—˝ ̵ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵“textbf–Parameter˝ ̵& ̵“textbf–Value˝ ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Number ̵of ̵source ̵positions ̵& ̵151 ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Source ̵spacing ̵ ̵& ̵2 ̵m ̵ ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵First ̵source ̵position ̵ ̵& ̵150.5““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Last ̵source ̵position ̵ ̵ ̵& ̵450.5 ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Number ̵of ̵receiver ̵positions ̵per ̵source ̵& ̵601 ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Receiver ̵spacing ̵& ̵1 ̵m ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵First ̵receiver ̵position ̵& ̵0 ̵m ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵Frequency ̵range ̵of ̵the ̵sweep ̵& ̵8-250 ̵Hz ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵“end–tabular˝ ̵ ̵“label–tab:5˙1˝“end–table˝“par%%%%%%%%%%%%%%% ̵Figure ̵5-3 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.9“textwidth]–Figures/Figure5˙3.png˝ ̵ ̵ ̵ ̵“caption–(a) ̵Location ̵of ̵the ̵testsite, ̵and ̵(b) ̵the ̵geometry ̵of ̵the ̵seismic ̵line. ̵The ̵red ̵stars ̵represent ̵active ̵sources, ̵the ̵blue ̵circles ̵represent ̵receivers, ̵and ̵the ̵orange ̵circles ̵represent ̵receivers ̵which ̵act ̵as ̵virtual ̵sources ̵for ̵applying ̵the ̵RISS.˝ ̵ ̵ ̵ ̵“label–fig:5˙3˝“end–figure˝%-------------------------------------------“section–Results˝The ̵primary ̵aim ̵of ̵this ̵study ̵is ̵to ̵examine ̵the ̵effectiveness ̵of ̵the ̵RISS ̵technique. ̵We ̵apply ̵the ̵technique ̵to ̵common-source ̵gathers ̵of ̵the ̵field ̵data, ̵as ̵introduced ̵in ̵Section ̵3. ̵The ̵data ̵processing ̵involves ̵several ̵steps. ̵Figure~“ref–fig:5˙4˝ ̵shows ̵a ̵flowchart ̵of ̵these ̵steps ̵for ̵the ̵RISS, ̵but ̵also ̵other ̵techniques ̵such ̵as ̵f-k ̵filtering ̵and ̵surgical ̵muting ̵for ̵surface-wave ̵suppression. ̵“par%%%%%%%%%%%%%%% ̵Figure ̵5-4 ̵%%%%%%%%%%%%%%“begin–figure˝[ht] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=“textwidth]–Figures/Figure5˙4.png˝ ̵ ̵ ̵ ̵“caption–Flowchart ̵for ̵applying ̵surface-wave ̵suppression. ̵RISS ̵represents ̵steps ̵for ̵the ̵Recursive ̵Interferometric ̵Surface-wave ̵Suppression, ̵the ̵RISS-muteR ̵is ̵the ̵same ̵as ̵RISS ̵but ̵using ̵muted ̵SI ̵results, ̵the ̵RISS-muteS ̵is ̵the ̵same ̵as ̵RISS ̵but ̵using ̵time-muted ̵reflection ̵dataset ̵as ̵input ̵for ̵SI, ̵SS-fk ̵represents ̵surface-wave ̵suppression ̵using ̵f-k ̵filtering, ̵SS-muteS ̵represents ̵surface-wave ̵suppression ̵using ̵time-muted ̵reflection ̵dataset.˝ ̵ ̵ ̵ ̵“label–fig:5˙4˝“end–figure˝As ̵shown ̵in ̵Figure~“ref–fig:5˙4˝, ̵first ̵we ̵apply ̵a ̵band-reject ̵filter ̵between ̵40 ̵Hz ̵and ̵90 ̵Hz ̵to ̵all ̵active ̵common-source ̵gathers. ̵The ̵main ̵idea ̵is ̵to ̵reject ̵frequencies ̵that ̵might ̵contain ̵reflections ̵but ̵not ̵surface ̵waves ̵so ̵that ̵the ̵SI ̵result ̵would ̵predominantly ̵contain ̵retrieved ̵surface ̵waves. ̵We ̵select ̵these ̵frequencies ̵based ̵on ̵the ̵power ̵spectrum ̵of ̵the ̵common-source ̵gathers. ̵We ̵then ̵apply ̵SI ̵as ̵described ̵in ̵section ̵2.1 ̵by ̵selecting ̵each ̵receiver ̵close ̵to ̵an ̵active ̵source ̵to ̵turn ̵it ̵into ̵a ̵virtual ̵source, ̵as ̵shown ̵by ̵the ̵orange ̵circles ̵in ̵Figure~“ref–fig:5˙3˝b. ̵“parNext, ̵we ̵adaptively ̵subtract ̵the ̵result ̵retrieved ̵by ̵SI ̵for ̵each ̵virtual ̵source ̵from ̵the ̵original ̵active-source ̵data ̵which ̵is ̵closest ̵to ̵the ̵virtual ̵source ̵to ̵suppress ̵the ̵surface ̵waves, ̵using ̵an ̵estimated ̵matching ̵filter ̵as ̵described ̵in ̵section ̵2.2. ̵“parIt ̵is ̵essential ̵to ̵determine ̵carefully ̵the ̵key ̵parameters ̵for ̵the ̵matching ̵filter: ̵time ̵window, ̵space ̵window, ̵and ̵filter ̵length. ̵We ̵choose ̵20 ̵traces ̵for ̵the ̵space ̵window ̵and ̵0.2 ̵s ̵for ̵time ̵time ̵window, ̵with ̵a ̵filter ̵length ̵of ̵0.05 ̵s. ̵We ̵apply ̵the ̵same ̵steps ̵for ̵all ̵virtual ̵sources. ̵Figure~“ref–fig:5˙5˝a ̵shows ̵an ̵example ̵of ̵the ̵original ̵common-source ̵gather ̵in ̵the ̵time ̵domain, ̵while ̵Figure~“ref–fig:5˙5˝b ̵shows ̵the ̵same ̵gather ̵after ̵RISS ̵with ̵one ̵iteration, ̵for ̵an ̵active ̵source ̵located ̵at ̵lateral ̵position ̵320.5 ̵m, ̵i.e., ̵after ̵the ̵SI ̵interferometric ̵suppression ̵as ̵applied ̵till ̵now. ̵“parAs ̵discussed ̵in ̵Section ̵2, ̵we ̵propose ̵to ̵apply ̵the ̵RISS ̵for ̵more ̵than ̵one ̵iteration ̵which ̵means ̵we ̵use ̵the ̵data ̵after ̵the ̵first ̵iteration ̵of ̵the ̵RISS ̵as ̵input ̵for ̵applying ̵SI. ̵Then, ̵we ̵repeat ̵all ̵steps ̵i.e., ̵band-reject ̵filtering, ̵SI, ̵and ̵adaptive ̵subtraction. ̵Note ̵that ̵these ̵steps ̵are ̵shown ̵as ̵"RISS" ̵in ̵the ̵flowchart ̵in ̵Figure~“ref–fig:5˙4˝.“parFigure~“ref–fig:5˙5˝c ̵shows ̵the ̵same ̵gather ̵as ̵in ̵Figures~“ref–fig:5˙5˝a ̵and ̵“ref–fig:5˙5˝b ̵but ̵after ̵the ̵second ̵iteration ̵of ̵the ̵RISS. ̵Comparing ̵these ̵three ̵results, ̵we ̵observe ̵that ̵a ̵large ̵part ̵of ̵the ̵surface-wave ̵energy ̵is ̵suppressed ̵in ̵Figure~“ref–fig:5˙5˝c, ̵as ̵shown ̵by ̵the ̵white ̵arrows. ̵Figures~“ref–fig:5˙5˝d-“ref–fig:5˙5˝f ̵show ̵the ̵f-k ̵spectra ̵of ̵the ̵common-source ̵gathers ̵in ̵Figures~“ref–fig:5˙5˝a-“ref–fig:5˙5˝c, ̵respectively. ̵The ̵surface-wave ̵energy ̵appears ̵as ̵linear ̵events ̵in ̵the ̵f-k ̵domain, ̵as ̵indicated ̵by ̵the ̵blue ̵arrows; ̵they ̵are ̵largely ̵suppressed ̵from ̵the ̵data ̵after ̵the ̵RISS ̵with ̵two ̵iterations, ̵as ̵can ̵be ̵observed ̵in ̵Figure~“ref–fig:5˙5˝f. ̵“parFigure~“ref–fig:5˙6˝ ̵shows ̵another ̵example ̵for ̵a ̵common-source ̵gather ̵for ̵a ̵source ̵located ̵at ̵a ̵lateral ̵position ̵of ̵430.5 ̵m, ̵where ̵we ̵also ̵observe ̵significant ̵suppression ̵of ̵the ̵surface-wave ̵energy ̵in ̵both ̵the ̵space-time ̵and ̵f-k ̵domain.“par%%%%%%%%%%%%%%% ̵Figure ̵5-5 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.95“textwidth]–Figures/Figure5˙5.png˝ ̵ ̵ ̵ ̵“caption–(a) ̵A ̵common-source ̵gather ̵for ̵a ̵source ̵located ̵at ̵lateral ̵position ̵320.5 ̵m, ̵(b) ̵same ̵common-source ̵gather ̵after ̵the ̵first ̵iteration ̵of ̵the ̵RISS, ̵(c) ̵same ̵common-source ̵gather ̵after ̵the ̵second ̵RISS ̵iteration; ̵(d), ̵(e) ̵and ̵(f) ̵same ̵as ̵(a), ̵(b), ̵and ̵(c), ̵respectively, ̵but ̵in ̵the ̵f-k ̵domain. ̵White ̵arrows ̵point ̵to ̵suppressed ̵surface ̵waves ̵in ̵the ̵space-time ̵domain ̵and ̵blue ̵arrows ̵point ̵to ̵surface-wave ̵energy ̵in ̵the ̵f-k ̵domain.˝ ̵ ̵ ̵ ̵“label–fig:5˙5˝“end–figure˝%%%%%%%%%%%%%%% ̵Figure ̵5-6 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.95“textwidth]–Figures/Figure5˙6.png˝ ̵ ̵ ̵ ̵“caption–Same ̵as ̵Figure~“ref–fig:5˙5˝ ̵but ̵for ̵a ̵source ̵at ̵a ̵lateral ̵position ̵430.5 ̵m.˝ ̵ ̵ ̵ ̵“label–fig:5˙6˝“end–figure˝“newpageAs ̵depicted ̵in ̵Figures~“ref–fig:5˙5˝c ̵and ̵“ref–fig:5˙6˝c, ̵we ̵successfully ̵suppress ̵the ̵surface ̵waves. ̵However, ̵it ̵appears ̵that ̵some ̵deeper ̵reflections ̵are ̵also ̵suppressed ̵in ̵the ̵process ̵(the ̵red ̵pointers ̵in ̵Figures ̵“ref–fig:5˙7˝a ̵and ̵“ref–fig:5˙7˝d). ̵This ̵shows ̵that ̵applying ̵a ̵simple ̵band-reject ̵filter ̵and ̵relying ̵on ̵having ̵less ̵sources ̵in ̵the ̵stationary-phase ̵regions ̵contributing ̵to ̵retrieval ̵of ̵reflections ̵might ̵not ̵guarantee ̵that ̵the ̵retrieved ̵reflection ̵energy ̵is ̵absent ̵or ̵even ̵weak. ̵To ̵preserve ̵these ̵reflections ̵in ̵the ̵original ̵common-source ̵gathers, ̵we ̵apply ̵a ̵bottom ̵muting ̵to ̵the ̵virtual-common-source ̵gathers ̵retrieved ̵from ̵SI ̵before ̵adaptive ̵subtraction, ̵which ̵we ̵label ̵as ̵"RISS-muteR" ̵in ̵the ̵flowchart ̵in ̵Figure~“ref–fig:5˙4˝. ̵Figures~“ref–fig:5˙7˝b ̵and ̵“ref–fig:5˙7˝e ̵show ̵the ̵common-source ̵gather ̵after ̵applying ̵the ̵RISS ̵using ̵the ̵muted ̵SI ̵results ̵for ̵two ̵active ̵sources ̵located ̵at ̵320.5 ̵m ̵and ̵430.5 ̵m, ̵respectively. ̵In ̵comparison ̵with ̵Figures~“ref–fig:5˙7˝a ̵and ̵“ref–fig:5˙7˝d, ̵which ̵show ̵the ̵same ̵common-source ̵gather ̵after ̵RISS, ̵we ̵observe ̵clearer ̵deeper ̵reflections ̵as ̵marked ̵by ̵the ̵red ̵arrows. ̵“parBy ̵examining ̵the ̵common-source ̵gathers, ̵we ̵observe ̵that ̵it ̵is ̵feasible ̵to ̵suppress ̵some ̵parts ̵of ̵the ̵surface ̵waves ̵through ̵surgical ̵muting, ̵which ̵is ̵a ̵common ̵approach. ̵Therefore, ̵prior ̵to ̵applying ̵the ̵RISS, ̵we ̵can ̵also ̵surgically ̵mute ̵the ̵prominent ̵surface ̵waves. ̵We ̵label ̵this ̵procedure ̵as ̵"RISS-muteS" ̵in ̵the ̵flowchart ̵in ̵Figure~“ref–fig:5˙4˝. ̵Figures~“ref–fig:5˙7˝c ̵and ̵“ref–fig:5˙7˝f ̵show ̵the ̵common-source ̵gather ̵after ̵the ̵RISS-muteS. ̵Although ̵we ̵enhance ̵the ̵resolution ̵of ̵some ̵reflections, ̵we ̵still ̵seems ̵to ̵have ̵some ̵strong ̵surface-wave ̵energy ̵in ̵comparison ̵with ̵the ̵results ̵in ̵Figures~“ref–fig:5˙7˝a ̵and ̵“ref–fig:5˙7˝d ̵as ̵highlighted ̵by ̵the ̵blue ̵ellipses. ̵This ̵observation ̵underscores ̵the ̵fact ̵that ̵by ̵suppressing ̵the ̵surface ̵waves ̵in ̵common-source ̵gathers ̵before ̵the ̵RISS, ̵it ̵becomes ̵challenging ̵to ̵retrieve ̵the ̵surface ̵waves ̵by ̵SI. ̵“par%%%%%%%%%%%%%%% ̵Figure ̵5-7 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.95“textwidth]–Figures/Figure5˙7.png˝ ̵ ̵ ̵ ̵“caption–(a) ̵A ̵common-source ̵gather ̵for ̵a ̵source ̵located ̵at ̵lateral ̵position ̵320.5 ̵m ̵after ̵the ̵RISS, ̵(b) ̵same ̵common-source ̵gather ̵after ̵the ̵RISS-muteR, ̵(c) ̵same ̵common-source ̵gather ̵after ̵the ̵RISS-muteS, ̵(d), ̵(e) ̵and ̵(f) ̵same ̵as ̵(a), ̵(b), ̵and ̵(c), ̵respectively, ̵but ̵for ̵a ̵source ̵located ̵at ̵a ̵lateral ̵position ̵of ̵430.5 ̵m. ̵Red ̵arrows ̵point ̵to ̵enhanced ̵reflections, ̵while ̵the ̵blue ̵ellipses ̵highlight ̵parts ̵of ̵the ̵surface ̵waves.˝ ̵ ̵ ̵ ̵“label–fig:5˙7˝“end–figure˝After ̵the ̵suppression ̵of ̵the ̵surface ̵waves, ̵we ̵apply ̵conventional ̵seismic ̵processing ̵to ̵obtain ̵preliminary ̵unmigrated ̵stacked ̵sections ̵for ̵a ̵better ̵comparison ̵between ̵the ̵result ̵of ̵the ̵RISS ̵and ̵the ̵conventional ̵suppression ̵techniques. ̵A ̵summary ̵of ̵these ̵steps ̵is ̵presented ̵in ̵Table~“ref–tab:5˙2˝. ̵We ̵first ̵apply ̵bandpass ̵filtering, ̵amplitude ̵correction ̵for ̵absorption ̵effects ̵and ̵geometrical ̵spreading, ̵and ̵then ̵automatic ̵gain ̵control ̵(AGC) ̵for ̵amplitude ̵balancing ̵for ̵visualization ̵purposes.“parNext, ̵we ̵sort ̵the ̵data ̵into ̵common-midpoint ̵(CMP) ̵gathers ̵(CMP ̵spacing ̵0.5 ̵m). ̵As ̵expected, ̵the ̵CMP ̵fold ̵increases ̵towards ̵the ̵center ̵of ̵the ̵line ̵where ̵better ̵illumination ̵is ̵achieved. ̵We ̵then ̵apply ̵normal ̵moveOut ̵(NMO) ̵correction ̵using ̵a ̵constant ̵velocity ̵of ̵350 ̵m/s, ̵and ̵finally, ̵we ̵stack ̵the ̵CMP ̵gathers. ̵The ̵constant ̵velocity ̵of ̵350 ̵m/s ̵is ̵selected ̵based ̵on ̵a ̵comparison ̵of ̵the ̵stacked ̵sections ̵with ̵different ̵velocity ̵values, ̵as ̵this ̵velocity ̵yields ̵better ̵results. ̵Figure~“ref–fig:5˙8˝ ̵shows ̵the ̵preliminary, ̵unmigrated ̵stacked ̵section ̵between ̵151.25 ̵m ̵and ̵450.25 ̵m ̵lateral ̵distances ̵for ̵the ̵five ̵approaches ̵for ̵surface-wave ̵suppression ̵as ̵illustrated ̵in ̵Figure~“ref–fig:5˙4˝. ̵Since ̵we ̵know ̵that ̵the ̵most ̵significant ̵influence ̵of ̵the ̵surface ̵waves ̵is ̵related ̵to ̵the ̵shallowest ̵part ̵of ̵the ̵subsurface, ̵we ̵focus ̵our ̵attention ̵to ̵these ̵parts, ̵specifically ̵400-800 ̵ms. ̵“parFigure~“ref–fig:5˙8˝a ̵shows ̵the ̵time ̵section ̵obtained ̵after ̵suppressing ̵the ̵surface ̵waves ̵using ̵surgical ̵muting, ̵as ̵indicated ̵by ̵SS-muteS ̵in ̵the ̵flowchart. ̵We ̵use ̵this ̵figure ̵as ̵a ̵reference ̵because ̵this ̵suppression ̵approach ̵is ̵standard ̵and ̵experienced ̵data ̵processors ̵generally ̵achieve ̵good ̵results. ̵This ̵result ̵is ̵comparable ̵with ̵results ̵from ̵other ̵studies, ̵e.g., ̵that ̵in ̵“textcite–Kruiver2017˝. ̵Figure~“ref–fig:5˙8˝b ̵shows ̵the ̵time ̵section ̵using ̵the ̵result ̵after ̵the ̵second ̵iteration ̵of ̵the ̵RISS. ̵In ̵comparison ̵with ̵the ̵reference ̵time ̵section, ̵it ̵is ̵obvious ̵that ̵some ̵of ̵the ̵expected ̵reflectors ̵are ̵suppressed ̵as ̵shown ̵by ̵the ̵light ̵blue, ̵red, ̵and ̵green ̵arrows. ̵“parTo ̵address ̵this ̵issue, ̵we ̵use ̵the ̵RISS-muteR ̵as ̵explained ̵above. ̵Figure~“ref–fig:5˙8˝c ̵illustrates ̵the ̵time ̵section ̵obtained ̵after ̵the ̵application ̵of ̵this ̵technique. ̵In ̵comparison ̵with ̵the ̵reference ̵in ̵Figure~“ref–fig:5˙8˝a, ̵we ̵retrieve ̵all ̵reflectors, ̵as ̵indicated ̵by ̵the ̵light ̵blue, ̵red, ̵and ̵green ̵arrows. ̵Moreover, ̵some ̵ ̵dome-like ̵structures ̵are ̵now ̵interpretable ̵as ̵highlighted ̵by ̵the ̵red ̵curves. ̵“parFigure~“ref–fig:5˙8˝d ̵shows ̵the ̵time ̵section ̵after ̵applying ̵RISS-muteS. ̵Comparing ̵this ̵image ̵to ̵the ̵images ̵in ̵Figures~“ref–fig:5˙8˝a,~“ref–fig:5˙8˝c, ̵we ̵see ̵that ̵the ̵lateral ̵continuity ̵of ̵the ̵reflectors ̵is ̵worse, ̵e.g., ̵inside ̵the ̵blue ̵ellipse, ̵while ̵the ̵general ̵character ̵on ̵the ̵left ̵part ̵of ̵the ̵image ̵has ̵become ̵lower ̵in ̵frequency, ̵which ̵might ̵point ̵to ̵the ̵presence ̵of ̵left-over ̵dominant ̵surface-wave ̵energy.Figure~“ref–fig:5˙8˝e ̵shows ̵the ̵time ̵section ̵obtained ̵after ̵suppression ̵of ̵the ̵surface ̵waves ̵by ̵f-k ̵filtering ̵(SS-fk), ̵which ̵is ̵another ̵common ̵surface-wave ̵suppression ̵technique. ̵Comparing ̵this ̵image ̵to ̵the ̵images ̵in ̵Figures~“ref–fig:5˙8˝a,~“ref–fig:5˙8˝c, ̵we ̵see ̵that ̵the ̵result ̵in ̵Figure~“ref–fig:5˙8˝e ̵is ̵generally ̵of ̵good ̵quality. ̵The ̵reflector ̵indicated ̵by ̵the ̵green ̵arrow ̵even ̵appears ̵laterally ̵more ̵continuous ̵than ̵in ̵Figures~“ref–fig:5˙8˝a,~“ref–fig:5˙8˝c. ̵However, ̵other ̵reflectors ̵are ̵less ̵clear, ̵e.g., ̵the ̵one ̵indicated ̵by ̵the ̵blue ̵arrow ̵or ̵those ̵on ̵in ̵the ̵left ̵part ̵of ̵the ̵image ̵earlier ̵than ̵600 ̵ms ̵(see ̵specifically ̵inside ̵the ̵blue ̵ellipse).“parBased ̵on ̵this ̵comparison, ̵it ̵appears ̵that ̵the ̵best ̵image ̵the ̵one ̵in ̵Figure~“ref–fig:5˙8˝c, ̵i.e., ̵after ̵application ̵of ̵RISS-muteR, ̵because ̵of ̵its ̵clarity ̵but ̵also ̵because ̵the ̵surface-wave ̵suppression ̵is ̵data-driven.“par“newpage“begin–table˝[htbp] ̵ ̵ ̵ ̵“caption–Summary ̵of ̵seismic ̵processing ̵steps.˝ ̵ ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵ ̵“small ̵ ̵ ̵ ̵“begin–tabular˝–—c—l—˝ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵“textbf–Step˝ ̵& ̵“textbf–Instruction˝ ̵““ ̵ ̵ ̵ ̵“hline ̵ ̵ ̵ ̵1 ̵& ̵Band-pass ̵filter ̵between ̵30-100 ̵Hz ̵““[0.1pt] ̵ ̵ ̵ ̵2 ̵& ̵Amplitude ̵corrections ̵““[0.1pt] ̵ ̵ ̵ ̵3 ̵& ̵Automatic ̵gain ̵control ̵(AGC) ̵ ̵““[0.1pt] ̵ ̵ ̵ ̵4 ̵& ̵Time ̵muting ̵ ̵ ̵““[0.1pt] ̵ ̵ ̵ ̵5 ̵& ̵Normal ̵moveout ̵(NMO) ̵correction ̵ ̵““[0.1pt] ̵ ̵ ̵ ̵6 ̵& ̵Common ̵midpoint/ensemble ̵stack ̵““[0.1pt] ̵ ̵ ̵ ̵“hline ̵ ̵“end–tabular˝ ̵ ̵“label–tab:5˙2˝“end–table˝%%%%%%%%%%%%%%% ̵Figure ̵5-8 ̵%%%%%%%%%%%%%%%“begin–figure˝[H]% ̵ ̵ ̵ ̵“centering% ̵ ̵ ̵ ̵“includegraphics[width=0.99“textwidth]–Figures/Figure5˙8.jpg˝% ̵ ̵ ̵ ̵“caption–Preliminary ̵unmigrated ̵stacked ̵section ̵using ̵a ̵constant ̵velocity ̵of ̵350 ̵m/s ̵using ̵raw ̵reflection ̵dataset ̵after ̵surgical ̵muting ̵of ̵the ̵surface ̵waves. ̵The ̵orange ̵arrows ̵point ̵to ̵strong ̵reflectors.˝% ̵ ̵ ̵ ̵“label–fig:5˙8˝%“end–figure˝%%%%%%%%%%%%%%% ̵Figure ̵5-8 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.9“textwidth]–Figures/Figure5˙8.jpg˝ ̵ ̵ ̵ ̵“caption–Preliminary, ̵unmigrated ̵stacked ̵section ̵using ̵a ̵constant ̵velocity ̵of ̵350 ̵m/s ̵zoomed ̵in ̵400-800 ̵ms, ̵(a) ̵using ̵the ̵reflection ̵dataset ̵after ̵surgical ̵muting ̵of ̵surface ̵waves ̵("SS-muteS"), ̵(b) ̵after ̵the ̵secon ̵interation ̵of ̵the ̵RISS, ̵(c) ̵after ̵the ̵RISS ̵using ̵the ̵muted ̵SI ̵results ̵(RISS-muteR), ̵(d) ̵using ̵the ̵data ̵as ̵in ̵(a) ̵but ̵after ̵the ̵RISS ̵("RISS-muteS"), ̵(e) ̵after ̵f-k ̵filtering ̵("SS-fk"). ̵The ̵colored ̵arrows ̵point ̵to ̵strong ̵reflectors ̵and ̵the ̵blue ̵ellipses ̵highlight ̵artefacts.˝ ̵ ̵ ̵ ̵“label–fig:5˙8˝“end–figure˝To ̵demonstrate ̵the ̵applicability ̵of ̵using ̵SI ̵for ̵surface-wave ̵suppression ̵in ̵imaging ̵the ̵shallow ̵subsurface, ̵we ̵further ̵investigate ̵its ̵use ̵for ̵the ̵Marchenko-based ̵isolation ̵(“cite–Wapenaar2021EMPLOYINGMETHOD˝; ̵“cite–VanIjsseldijk2023ExtractingIsolation˝). ̵The ̵Marchenko ̵method ̵can ̵estimate ̵Green's ̵functions ̵between ̵the ̵Earth's ̵surface ̵and ̵arbitrary ̵locations ̵in ̵the ̵subsurface. ̵These ̵Green's ̵functions ̵are ̵used ̵to ̵redatum ̵wavefields ̵to ̵a ̵deeper ̵level ̵in ̵the ̵subsurface ̵(“cite–Slob2013SeismicEquations˝; ̵“cite–Wapenaar2014MarchenkoImaging˝). ̵The ̵Marchenko ̵method ̵can ̵also ̵be ̵used ̵to ̵isolate ̵the ̵response ̵of ̵a ̵specific ̵layer ̵or ̵package ̵of ̵layers, ̵free ̵from ̵the ̵influence ̵of ̵the ̵overburden ̵and ̵underburden. ̵The ̵complete ̵derivation ̵of ̵the ̵Marchenko-based ̵isolation ̵is ̵beyond ̵the ̵scope ̵of ̵this ̵paper; ̵however, ̵a ̵detailed ̵description ̵can ̵be ̵found ̵in ̵“textcite–VanIjsseldijk2023ExtractingIsolation˝, ̵“textcite–IJsseldijkTrollb˝. ̵Moreover, ̵a ̵complete ̵description ̵of ̵the ̵application ̵of ̵the ̵Marchenko-based ̵isolation ̵on ̵the ̵field ̵dataset ̵used ̵in ̵this ̵study ̵can ̵be ̵found ̵in ̵“textcite–FaezehThesis˝. ̵“parWe ̵use ̵the ̵Marchenko-based ̵isolation ̵to ̵eliminate ̵the ̵overburden ̵and ̵the ̵underburden, ̵and ̵isolate ̵the ̵reflection ̵from ̵the ̵target ̵layer ̵between ̵30 ̵m ̵and ̵270 ̵m ̵using ̵the ̵data ̵after ̵surgical ̵muting ̵of ̵surface ̵waves ̵and ̵the ̵data ̵after ̵RISS-muteR. ̵Figure~“ref–fig:5˙9˝a ̵shows ̵the ̵stacked ̵section ̵using ̵the ̵regular ̵reflection ̵response ̵after ̵suppression ̵of ̵the ̵surface ̵waves ̵using ̵surgical ̵muting, ̵while ̵Figures~“ref–fig:5˙9˝b ̵and~“ref–fig:5˙9˝c ̵show ̵the ̵stacked ̵section ̵using ̵the ̵reflection ̵response ̵after ̵Marchenko-based ̵isolation ̵for ̵overburden ̵and ̵underburden ̵removal, ̵using ̵surgical ̵muting ̵for ̵surface-wave ̵suppression ̵and ̵RISS-muteR, ̵respectively. ̵“parA ̵comparison ̵of ̵these ̵stacked ̵section ̵in ̵Figure~“ref–fig:5˙9˝ ̵suggests ̵the ̵potential ̵elimination ̵of ̵the ̵internal ̵multiples ̵originating ̵from ̵the ̵overburden, ̵down ̵to ̵30 ̵m ̵using ̵the ̵Marchenko-based ̵isolation. ̵The ̵shallow ̵reflectors ̵appear ̵clearer ̵and ̵more ̵continuous, ̵as ̵indicated ̵by ̵the ̵color-coded ̵arrows. ̵But ̵what ̵we ̵want ̵to ̵draw ̵attention ̵to ̵is ̵the ̵effect ̵of ̵surface-wave ̵suppression ̵on ̵these ̵results. ̵We ̵observe ̵enhanced ̵reflections ̵with ̵a ̵ ̵greater ̵clarity ̵of ̵the ̵stacked ̵sections ̵after ̵RISS-muteR ̵(Figure~“ref–fig:5˙9˝c), ̵as ̵exemplified ̵by ̵the ̵events ̵indicated ̵by ̵the ̵white ̵and ̵green ̵arrows. ̵Moreover, ̵there ̵are ̵fewer ̵artefacts, ̵likely ̵from ̵surface ̵waves, ̵as ̵indicated ̵by ̵the ̵black ̵ellipse. ̵“par%%%%%%%%%%%%%% ̵Figure ̵5-9 ̵%%%%%%%%%%%%%%“begin–figure˝[H] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=“textwidth]–Figures/Figure5˙9.png˝ ̵ ̵ ̵ ̵“caption–Stacked ̵sections, ̵zoomed ̵in ̵between ̵0.6 ̵s ̵and ̵1.0 ̵s, ̵obtained ̵using ̵(a) ̵the ̵regular ̵reflection ̵response ̵after ̵surgical ̵muting ̵of ̵the ̵surface ̵waves, ̵(b) ̵the ̵reflection ̵response ̵after ̵Marchenko-based ̵isolation ̵for ̵overburden ̵and ̵underburden ̵removal ̵after ̵surgical ̵muting ̵of ̵the ̵surface ̵waves, ̵and ̵(c) ̵similar ̵to ̵(b) ̵but ̵using ̵data ̵after ̵surface-wave ̵suppression ̵with ̵RISS-muteR ̵instead ̵of ̵surgical ̵muting. ̵The ̵colour-coded ̵arrows ̵indicate ̵reflectors. ̵The ̵black ̵ellipse ̵highlights ̵potential ̵artefacts ̵from ̵the ̵surface ̵waves ̵that ̵are ̵suppressed ̵in ̵(c).˝ ̵ ̵ ̵ ̵“label–fig:5˙9˝“end–figure˝%-------------------------------------------“section–Discussion˝We ̵presented ̵a ̵comparison ̵of ̵different ̵approaches ̵for ̵surface-wave ̵suppression ̵applied ̵to ̵the ̵land ̵seismic ̵dataset ̵acquired ̵in ̵Scheemda, ̵Groningen ̵province: ̵surgical ̵muting, ̵f-k ̵filtering, ̵the ̵RISS, ̵the ̵RISS-muteR, ̵and ̵the ̵RISS-muteS. ̵“parFrom ̵a ̵comparison ̵of ̵the ̵unmigrated ̵time ̵sections, ̵we ̵found ̵that ̵surgical ̵muting ̵seems ̵convenient ̵for ̵removing ̵surface ̵waves ̵in ̵our ̵dataset. ̵However, ̵this ̵method ̵does ̵not ̵adequately ̵remove ̵the ̵surface-wave ̵energy, ̵as ̵surface ̵waves ̵overlap ̵with ̵useful ̵reflections ̵and ̵scattered ̵arrivals. ̵Additionally, ̵weak ̵reflections ̵and ̵scattered ̵arrivals ̵covered ̵by ̵surface ̵waves ̵might ̵also ̵be ̵muted, ̵as ̵demonstrated ̵earlier ̵ ̵by ̵“textcite–KONSTANTAKI201528˝.The ̵other ̵common ̵technique ̵for ̵surface-wave ̵suppression ̵-f-k ̵filtering- ̵could ̵cause ̵more ̵artefacts, ̵as ̵it ̵is ̵challenging ̵to ̵define ̵the ̵correct ̵parameters ̵for ̵frequency ̵and ̵wavenumber ̵windows ̵for ̵all ̵common-source ̵gathers. ̵In ̵contrast, ̵the ̵RISS ̵technique ̵can ̵effectively ̵suppress ̵surface ̵waves ̵without ̵prior ̵information, ̵as ̵it ̵is ̵purely ̵data-driven. ̵“parBy ̵comparing ̵three ̵approaches ̵for ̵the ̵RISS, ̵we ̵found ̵that ̵the ̵RISS ̵using ̵muted ̵deeper ̵reflections ̵after ̵retrieving ̵the ̵surface ̵waves ̵with ̵SI ̵yields ̵the ̵best ̵results. ̵RISS ̵with ̵two ̵iterations ̵was ̵sufficient ̵to ̵achieve ̵the ̵desired ̵results ̵for ̵our ̵dataset. ̵However, ̵for ̵more ̵than ̵two ̵iterations ̵can ̵be ̵needed ̵for ̵other ̵datasets.“parFor ̵implementation ̵of ̵the ̵RISS, ̵we ̵must ̵determine ̵the ̵time, ̵spacing ̵window, ̵and ̵the ̵filter ̵length. ̵We ̵determined ̵these ̵parameters ̵carefully ̵by ̵examining ̵different ̵values ̵for ̵them. ̵“textcite–Balestrini2020ImprovedSuppression˝ ̵demonstrated ̵that ̵changing ̵the ̵time ̵and ̵the ̵space ̵window ̵size ̵makes ̵no ̵significant ̵differences. ̵They ̵also ̵observed ̵that ̵increasing ̵the ̵filter ̵length ̵produces ̵undesirable ̵artefacts ̵at ̵earlier ̵times. ̵Defining ̵a ̵proper ̵filter ̵length ̵can ̵be ̵indeed ̵crucial ̵for ̵different ̵datasets. ̵“parA ̵comparison ̵of ̵the ̵surface-wave ̵dispersion ̵curves ̵obtained ̵from ̵the ̵results ̵of ̵SI ̵with ̵active ̵common-source ̵gathers ̵shows ̵that ̵the ̵kinematics ̵of ̵the ̵higher ̵modes ̵of ̵surface ̵waves ̵cannot ̵be ̵accurately ̵retrieved ̵using ̵SI, ̵as ̵depicted ̵in ̵Figure~“ref–fig:5˙10˝. ̵Consequently, ̵effective ̵suppression ̵of ̵these ̵higher ̵modes ̵poses ̵a ̵significant ̵challenge. ̵As ̵highlighted ̵by ̵“textcite–Halliday2008˝, ̵modal ̵separation ̵is ̵possibly ̵a ̵crucial ̵step ̵before ̵applying ̵SI ̵to ̵ensure ̵the ̵correct ̵kinematic ̵retrieval ̵and ̵consecutive ̵suppression ̵of ̵the ̵ ̵higher ̵modes ̵with ̵minimal ̵error. ̵On ̵the ̵other ̵hand, ̵the ̵RISS ̵can ̵help ̵suppress ̵the ̵higher ̵modes, ̵as ̵we ̵have ̵demonstrated ̵here. ̵This ̵happens ̵as, ̵during ̵each ̵iteration, ̵the ̵strongest ̵present ̵surface-wave ̵mode ̵is ̵retrieved ̵and ̵adaptively ̵suppressed, ̵which ̵could ̵effectively ̵be ̵seen ̵as ̵step-wise ̵modal ̵separation. ̵“parAs ̵described ̵in ̵the ̵methodology ̵section, ̵we ̵utilised ̵the ̵field ̵dataset ̵with ̵surface ̵waves ̵for ̵each ̵active ̵source ̵as ̵a ̵reference, ̵and ̵for ̵the ̵virtual ̵sources, ̵the ̵surface ̵waves ̵were ̵retrieved ̵using ̵SI. ̵In ̵our ̵study, ̵we ̵created ̵at ̵receiver ̵locations ̵as ̵the ̵virtual ̵sources, ̵which ̵are ̵positioned ̵at ̵0.5 ̵m ̵from ̵the ̵active ̵sources; ̵this ̵provides ̵an ̵appropriate ̵dataset ̵for ̵applying ̵this ̵technique. ̵However, ̵using ̵virtual ̵sources ̵located ̵at ̵greater ̵distances ̵from ̵the ̵active ̵sources ̵may ̵result ̵in ̵incorrect ̵estimations ̵of ̵the ̵surface ̵waves ̵using ̵SI. ̵Therefore, ̵we ̵recommend ̵using ̵sources ̵and ̵receivers ̵in ̵close ̵proximity ̵of ̵each ̵other ̵to ̵ensure ̵the ̵accuracy ̵of ̵RISS. ̵“parFinally, ̵besides ̵the ̵application ̵of ̵SI ̵for ̵surface-wave ̵suppression, ̵it ̵is ̵also ̵possible ̵to ̵utilise ̵SI ̵for ̵surface-wave ̵retrieval ̵at ̵each ̵receiver, ̵which ̵can ̵be ̵achieved ̵by ̵applying ̵an ̵extra ̵step- ̵a ̵subtraction ̵of ̵the ̵result ̵of ̵RISS ̵from ̵the ̵original ̵dataset, ̵resulting ̵in ̵a ̵virtual ̵source ̵gather ̵with ̵only ̵surface ̵waves ̵at ̵each ̵receiver ̵position. ̵This ̵approach ̵would ̵be ̵beneficial ̵in ̵cases ̵where ̵more ̵sources ̵are ̵needed, ̵such ̵as ̵for ̵full-waveform ̵inversion ̵of ̵surface ̵waves.“par%%%%%%%%%%%%%%% ̵Figure ̵5-10 ̵%%%%%%%%%%%%%%“begin–figure˝[h] ̵ ̵ ̵ ̵“centering ̵ ̵ ̵ ̵“includegraphics[width=0.8“textwidth]–Figures/Figure5˙10.png˝ ̵ ̵ ̵ ̵“caption–Dispersion ̵curves ̵(a) ̵for ̵an ̵active ̵source ̵at ̵320.5 ̵m, ̵(b) ̵the ̵virtual ̵source ̵at ̵320 ̵m, ̵(c) ̵same ̵as ̵(a) ̵but ̵for ̵the ̵active ̵source ̵at ̵430.5 ̵m, ̵and ̵(d) ̵same ̵as ̵(b) ̵but ̵for ̵the ̵virtual ̵source ̵at ̵430 ̵m. ̵The ̵black ̵arrows ̵point ̵to ̵higher ̵modes ̵of ̵surface ̵waves.˝ ̵ ̵ ̵ ̵“label–fig:5˙10˝“end–figure˝%-------------------------------------------“section–Conclusion˝We ̵proposed ̵recursive ̵application ̵of ̵seismic ̵interferometry ̵(SI) ̵for ̵surface-wave ̵suppression. ̵We ̵showcased ̵our ̵technique ̵using ̵a ̵2D ̵reflection ̵dataset ̵acquired ̵in ̵Scheemda, ̵Groningen ̵province, ̵the ̵Netherlands. ̵We ̵applied ̵SI ̵to ̵retrieve ̵dominant ̵surface ̵waves ̵between ̵receivers ̵while ̵minimising ̵the ̵retrieved ̵reflection ̵energy. ̵The ̵retrieved ̵dominant ̵surface ̵waves ̵were ̵then ̵adaptively ̵subtracted ̵from ̵the ̵original ̵data. ̵We ̵showed ̵that ̵applying ̵these ̵two ̵steps ̵two ̵times, ̵i.e., ̵recursively, ̵resulted ̵in ̵a ̵fully ̵data-driven ̵effective ̵suppression ̵of ̵the ̵surface ̵waves. ̵“parWe ̵compared ̵stacked ̵sections ̵obtained ̵through ̵the ̵recursive ̵interferometric ̵surface ̵wave ̵suppression ̵with ̵stacked ̵sections ̵where ̵the ̵surface ̵waves ̵were ̵suppressed ̵using ̵f-k ̵filtering ̵and ̵surgical ̵muting. ̵We ̵found ̵that ̵the ̵obtained ̵time ̵section ̵after ̵the ̵second ̵iteration ̵of ̵recursive ̵interferometric ̵surface-wave ̵suppression ̵yielded ̵better ̵results ̵in ̵terms ̵of ̵clearer ̵and ̵more ̵continuous ̵reflections, ̵especially ̵when ̵the ̵SI ̵result ̵was ̵used ̵in ̵which ̵the ̵retrieved ̵reflection ̵energy ̵was ̵minimised ̵by ̵bottom ̵muting ̵before ̵SI. ̵This ̵technique ̵can ̵be ̵effective ̵for ̵enhancing ̵the ̵resolution ̵of ̵the ̵seismic ̵reflection ̵events ̵for ̵subsurface ̵investigations ̵or ̵alternatively ̵for ̵obtaining ̵surface ̵waves ̵free ̵from ̵body-wave ̵reflections. ̵“par%-------------------------------------------“section–Acknowledgement˝We ̵acknowledge ̵the ̵use ̵of ̵computational ̵resources ̵provided ̵by ̵the ̵DelftBlue ̵supercomputer ̵at ̵the ̵Delft ̵High ̵Performance ̵Computing ̵Centre ̵(“url–https://www.tudelft.nl/dhpc˝) ̵and ̵thank ̵Seismic ̵Mechatronics ̵for ̵allowing ̵us ̵to ̵use ̵their ̵vibrator ̵source ̵during ̵fieldwork. ̵(“url–https://seismic-mechatronics.com/˝) ̵“par“noindent ̵This ̵research ̵is ̵funded ̵by ̵NWO ̵Science ̵domain ̵(NWO-ENW), ̵project ̵DEEP.NL.2018.048 ̵and ̵the ̵reflection ̵dataset ̵was ̵acquired ̵by ̵funding ̵from ̵the ̵European ̵Research ̵Council ̵(ERC) ̵under ̵the ̵European ̵Union’s ̵Horizon ̵2020 ̵research ̵and ̵innovation ̵programme ̵(Grant ̵Agreement ̵No. ̵742703) ̵and ̵ ̵OYO ̵Corporation, ̵Japan: ̵OYO-TUD ̵Research ̵Collaboration ̵(Project ̵code ̵C25B74).%-------------------------------------------“section–Data ̵availability˝The ̵field ̵reflection ̵dataset ̵used ̵in ̵this ̵study ̵is ̵available ̵in ̵the ̵4TU.ResearchData ̵repository ̵at ̵“url–https://doi.org/10.4121/a8553b7e-82ae-4e9b-bc54-2a6b9ca6063c˝.““The ̵codes ̵are ̵also ̵accessible ̵in ̵the ̵4TU.ResearchData ̵repository ̵at ̵“url–https://doi.org/10.4121/6271c7d3-f931-49e9-b2b0-2d05eef7d3ae˝.%-------------------------------------------% ̵References%-------------------------------------------% ̵Print ̵bibliography“printbibliography“end–document˝